OpenStudy (shamil98):

\[\text{Find} \frac{ \delta z }{ \delta x }\text{and} \frac{ \delta z }{ \delta y } \text{ if z is defined implicitly as a function of x and y by the equation}\] \[x^3 + y^3 + z^3 + 6xyz = 1\]

OpenStudy (shamil98):

To differentiate implicitly dz/dx with respect to x i would treat y as a constant right?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

then solve for dz/dx

OpenStudy (usukidoll):



OpenStudy (primeralph):

Notation?

OpenStudy (usukidoll):

partial derivatives...

OpenStudy (usukidoll):

wait what I'm I doing here? I should be chillin... XD

OpenStudy (shamil98):

\[3x^2 + 3z^2 \frac{ \delta z }{ \delta x } + 6yz + 6xy \frac{ \delta z }{ \delta x } = 0\] \[\frac{ \delta z }{ \delta x } (3z^2 + 6xy) = 3(-x^2 -2 yz)\] \[\frac{ \delta z }{ \delta x } = \frac{ 3(-x^2 - 2yz) }{ 3(z^2 + 2xy) }\] \[ = \frac{ -x^2 - 2yz }{ z^2 + 2xy }\]

OpenStudy (primeralph):

Are those partials?

OpenStudy (shamil98):

yeah

OpenStudy (usukidoll):

what the? can't be that long

OpenStudy (primeralph):